A Suppose X and Y are random variables that only take on the values 0, 1, and 2. (That is, for both of their probability mass functions, p(z) = 0 for xメ0.1.2.) Suppose E(X-Ely and E X2 EY]. Prove that EEY (Write your answer in complete sentences, as this requires a proof.)
Answer #1

Let p(x) and q(y) be the pmf of x and y respectively.
ElX) = ElY/> Σ.rpl.r) = Σ lwly)
p(1) + 2p(2) = q(1) + 2q(2) _(I)

E[X^{2}] = E[Y^{2}] \Rightarrow \sum_{0}^{2}x^{2}p(x) = \sum_{0}^{2}y^{2}q(y)
p(1) + 4p(2) = q(1) + 4q(2) _(II)
Multiplying equation (I) by 2 we get
2p(1) + 4p(2) = 2q(1) + 4q(2) _(III)
Subtracting equation (III) from equation (I) we get
p(1) = q(1)
Subtracting equation (I) from equation (II) we get
2p(2) = 2q(2)
p(2) = q(2)
Since p(0) and q(0) = 0 \forall x,y\neq 0,1,2
and  p(1) = q(1) , p(2) = q(2)\Rightarrow p(0) = q(0)
Since the distribution of p(x) and q(y) are equal:
E[X3] = E[Y3]